OpenStudy (anonymous):

How to resolve it into partial fractions: \[\frac{1}{(u^2-1)^2}\]

OpenStudy (anonymous):

note that\[\frac{1}{(u^2-1)^2}=\frac{1}{(u-1)^2(u+1)^2}\]

OpenStudy (anonymous):

Yes... But how to continue?

OpenStudy (anonymous):

substitute 1 by 1/4((u+1)-(u-1))^2

OpenStudy (anonymous):

open the bracket cancel the terms and then u will get it simplified...

OpenStudy (anonymous):

@RolyPoly

OpenStudy (anonymous):

How can you think of this term: 1/4((u+1)-(u-1))^2 ?

OpenStudy (anonymous):

mathematical logic u just have to substitiue it in such a way that the denominator cancels out and u get ur task done... u will have to think it .. no formula for it..

OpenStudy (anonymous):

Hmm.. even though I don't get this ''mathematical logic'', but thanks still! I think I can work that partial fraction out from your hint!